Prove/ Disprove :
If A and B are any two events, then it is always true that
P (A ∪ B) ≤ P (A) + P (B)
Answer #1
P (A ∪ B) ≤ P (A) + P (B).
This is true

we know that
Probability is always non-negative(>=0)
and P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
since P(A ∩ B) is >= 0
so, P(A ∪ B) ≤ P(A) + P(B), because P(A ∩ B) is >= 0